OpenStudy (anonymous):

What point in the feasible region maximizes C for the objective function C=3x+2y?

OpenStudy (anonymous):

Constraints: x+2y<=10 3x-y<=9 x>=0, y >=0

OpenStudy (anonymous):

My last question of the day by the way!

OpenStudy (anonymous):

Im not sure about this one : eq 1 .3x - y = 9 eq 2. x + 2y = 10 eq x (2) : 6x - 2y = 18 + x + 2y = 10 7 x = 18 x= 18/7 y = 3x - 9 = 3 ( 18/7 ) - 9 = -9/7 C = 3 ( 18/7 ) + 2 ( -9/7 ) =36/7 not correct i guess..

OpenStudy (anonymous):

wait wait wait : error spotted ( sorry im a lil tired )..

OpenStudy (anonymous):

You've been a big help today, so if you need to sleep that's alright. :)

OpenStudy (anonymous):

Im not sure about this one : eq 1 .3x - y = 9 eq 2. x + 2y = 10 eq x (2) : 6x - 2y = 18 + x + 2y = 10 7 x = 28 x= 4 y = 9 + 3 x = 9 + 3 ( 4 ) = 21 => ( 4 , 21 )

OpenStudy (anonymous):

@hartnn .. check my answer :)